OpenStudy (anonymous):

<1 and <2 are supplementary angles. m <1 is 4y + 7 and m <2 is 9y + 4. Find m <2. Thanks

OpenStudy (anonymous):

M a1 + m a2 = 180 = def. of supplementary angles A1= 4y + 7 A2 = 9y + 4 a1 + a2 = 180 4y + 7 + 9y + 4 =180 13y +16 = 180 13y + 16 -16 = 180 - 16 now solve for y and use to find A1= and A2=?